Cycle Time, Velocity, Product Costing Mulhall, Inc., has a JIT system in place. Each manufacturing cell is dedicated to the production of a single product or major subassembly. One cell, dedicated to the production of mopeds, has four operations: machining, finishing, assembly, and qualifying (testing). The machining process is automated, using computers. In this process, the model’s frame and engine are constructed. In finishing, the frame is sandblasted, buffed, and painted. In assembly, the frame and engine are assembled. Finally, each model is tested to ensure operational capability.

For the coming year, the moped cell has the following budgeted costs and cell time (both at theoretical capacity):

Budgeted conversion costs $5,400,000

Budgeted materials $18,647,000

Cell time 36,000

Theoretical output 27,000 models

During the year, the following actual results were obtained:

Actual conversion costs $5,400,000

Actual materials $4,012,000

Actual cell time 36,000 hours

Actual output 22,500 models

Required: 1. Compute the velocity (number of models per hour) that the cell can theoretically achieve. If required, round your answer to two decimal places.

??? model per hour

Now, compute the theoretical cycle time (number of minutes per model) that it takes to produce one model. In your calculations, do not round the theoretical velocity.

???minutes per model

2. Compute the actual velocity and the actual cycle time. If required, round your actual velocity answer to three decimal places.

Actual velocity ???model per hour

Actual cycle time ???minutes per model

3. Compute MCE. If required, round your answer to two decimal places. Comment on the efficiency of the operation. ??

4. Compute the budgeted conversion cost per minute. If required, round your answer to the nearest cent. $ per minute

Using this rate, compute the conversion cost per model if theoretical output is achieved. $ per model

Using this measure, compute the conversion cost per model for actual output. $ per model

Does this product costing approach provide an incentive for the cell manager to reduce cycle time?

Yes or No?
Answer #1

1. Theoretical velocity = Theoretical output / Cell time

= 27,000/36,000 = 0.75 model per hour

Theoretical cycle time = No of minutes in a hour / Theoretical velocity

= 60/0.75 = 80 minutes per model

2. Actual velocity = Actual output / Cell time

= 22,500/36,000 = 0.625 model per hour

Actual cycle time = No of minutes in a hour / Actual velocity

= 60/0.625 = 96 minutes per model

3. MCE = Theoretical cycle time / Actual cycle time

= 80/96 = 0.83

The efficiency of the operation is very high as actual cycle time is greater than theoretical cycle time.

4. Budgeted conversion cost per minute = Budgeted conversion cost / (Cell time * No of minutes in a hour)

= $5,400,000 / (36,000 * 60) = $2.5 per minute

Theoretical conversion cost per model = Budgeted conversion cost per minute * Theoretical cycle time

= $2.5 × 80 = $200

Actual conversion cost per model = Budgeted conversion cost per minute * Actual cycle time

= $2.5 × 96 = $240

Yes. By reducing cycle time, the cost per unit can be reduced.

Reduction is = $240 – $200 = $40 per model